OpenStudy (anonymous):

The sales price of a car is $12,590, which is 20% off the original price. What is the original price? please explain it to me

OpenStudy (nowhereman):

Let the original price be denoted by x, then write down the function x ↦ $12590, that takes original prices and returns the sales price. Then you only need to invert it.

OpenStudy (anonymous):

ok and how would i do that math is my weakest subject its like disease that i cnt fight

OpenStudy (anonymous):

i didn't quite get the q so taking ur q as 12590 being the sale price with a 20% discount, x-0.2x=0.8x=12590 x=12590/0.8=15737.5

OpenStudy (anonymous):

ok, let me start from the beginning

OpenStudy (anonymous):

sales price (100% minus the 20% sales = 80% of the original)= $12590 original price (100%) = \[sales price \div 80 * 100\] = 12590/80*100 = 15737.50$

OpenStudy (anonymous):

this works like this: you know the price after a 20% off, which means the sales price you know is 80% of the original price. So to find the original price, you have to divide the sales price by 80 (to find how much $ is for each %), and then multiply by 100 so you get the total 100% price

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea but quick question

OpenStudy (amistre64):

12590 = original price - original price(%) 12590 = P - P(20%) change 20% to its decimal form (.20) 12590 = P(1 - .20) P = 12590/(1-.20)

OpenStudy (anonymous):

amistres64 how did you get one for this problem when you said (1-.20)

OpenStudy (amistre64):

It is a technique used to "factor out" a variable. Lets use P =10 and the % =3 as an easy example: just replace them with numbers: 10(1) - 10(3) we can easily find the value of this; 10 - 30 = -20 right? We can also solve this by "factoring out" the thing they have in common and turning it into a multiplication problem like this: 10 (1-3) which is 10(-2) = -20, same answer. Just remember that there is an "invisible" or "implied" (1) next to all variables and numbers; why? because anything times (1) equals itself. Does this make sense?

OpenStudy (amistre64):

In other words: P(1) - P(.20) factors out to be: P(1-.20)

OpenStudy (anonymous):

yes thank you ... you helped a lot

OpenStudy (anonymous):

so when you say 12590/(1-.20) are you telling me to divide or multiply before i subtract

OpenStudy (amistre64):

12590 divided by (1-.20)

OpenStudy (anonymous):

ok thank you so so so so much your a lifesaver